# The underlying structures of self-attention: symmetry, directionality, and emergent dynamics in Transformer training

### v1

@@ 4.1. Evolution of symmetric and directional structures during learning
Figure 3: a) Evolution of symmetry score during training. Shown are the median and the interquartile range. Models were trained on the Wikipedia dataset [Foundation, 2022]. Encoder-only and decoder-only models are color-coded in blue and purple, respectively (see legend). b) Same as in panel a for the median directionality score. c) Evolution of the median symmetry score across layers of the encoder-only (left) and decoder-only (right) models. Each layer is color-coded as shown on the legend. d) Same as panel c for the median directionality score.
To test the applicability of our result, we first train 12-layer transformer models in both encoder and decoder modes and quantify the median symmetry and directionality scores across epochs. At initialization, the symmetry and directionality score of the matrix $\mathbf{W}_{qk}$ at any layer is zero (see Definition 3.1 and Definition 3.2 and related Appendix S1.5 and S1.6). The incremental update of $\mathbf{W}_{qk}$ we described in the previous sections predicts that decoder-only models develop high-norm columns incrementally during training (see Theorem 2.3). Likewise, as symmetric weight updates are added to $\mathbf{W}_{qk}$ in encoder-only models, Theorem 2.4 predicts that symmetric structures emerge incrementally during training.
@@ 4.2. Enforcing symmetry at initialization improves the training of encoder-only models
Table 1: The final loss at the end of training and the speed-up for the 4 and 12-layer models trained on the Jigsaw dataset [cjadams et al., 2017], Wikipedia [Foundation, 2022], and Red Pajama [Computer, 2023], with and without symmetry initialization (see Appendix S2.1). Speed-up (%) is calculated by subtracting the epoch at which the symmetrically initialized model reaches the non-symmetric model’s final loss from the total number of epochs, and then dividing by the total number of epochs. For example, a 50% speed-up means that the model with symmetric initialization achieves the final loss of the non-symmetric model in half the number of training epochs.
@@
We train 4-layer and 12-layer encoder-only models, comparing two initialization strategies: initialize the self-attention matrices independently versus initializing the $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ in each self-attention layer to ensure that $\mathbf{W}_{qk}$ is symmetric (see Appendix S2). We report the results of our experiments in Table 1. We observe that enforcing symmetry at initialization leads to lower loss values at the end of training for most of the models. Importantly, symmetric initialization significantly accelerates convergence, reaching the final loss value faster than those with random initialization (up to ${}75\%$ faster for 4-layer models and ${}35\%$ faster for 12-layer models, see also Figure S5a). Moreover, we observe that self-attention matrices initialized symmetrically lose symmetry during training but converge to higher symmetry levels than random initialization (Figure S5b) This symmetric initialization decreases the gap in symmetric scores between layers compared to random initialization (Figure S5c). These results highlight that embedding symmetry as an inductive bias across all Transformer layers can enhance training efficiency and model performance.
@@ Acknowledgments
MS was supported by an ETH Postdoctoral fellowship (reference nr. 23-2 FEL-042). PVA was supported by a UZH Postdoctoral fellowship (reference nr. K-76207-01-01). TS and PS were supported by DIZH Fellowship “Stability of self-organizing net fragments as inductive bias for next-generation deep learning” from the ZHAW. We thank the Grewe lab, Alexander Nedergaard, Thomas Hofmann, and the Data Analytics lab for their helpful comments and insightful discussions.
@@
(Autoregressive and bidirectional objectives) Let $U=\{t_{1},\dots,t_{N}\}$ a sequence of tokens. The joint probability of $U$ is factorized autoregressively as follows,
@@
During autoregressive training, a model with set of parameters $\mathcal{W}$ is optimized to minimize the following negative log-likelihood
$\mathcal{L}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p(t_{i}\,|\,\{t_{j}\}:j<i\,;\mathcal{W})\,,$ (S7)
where the conditional probabilities are modeled with learnable parameters $\mathcal{W}$. The join probability of $U$ can also be factorized bidirectionally as follows,
$\Pr[U]=\Pr[t_{1},\dots,t_{N}]=\Pi_{i=1}^{N}\Pr[t_{i}|t_{1},\dots,t_{i-1},t_{i+1},\dots,t_{N}]\,.$ (S8)
During bidirectional training, a model with set of parameters $\mathcal{W}$ is optimized to minimize the following negative log-likelihood
$\mathcal{L}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p(t_{i}\,|\,\{t_{j}\}:j\neq i\,;\mathcal{W})\,.$ (S9)
@@
$\mathcal{L}(U;\mathcal{W})=-\sum_{i\in M}\log p(t_{i}\,|\,\{t_{j}\}:j\notin M\,;\mathcal{W})\,.$ (S10)
S1.2. Proof of Proposition 2.1 and related remarks
Proof.
Let $\mathbf{A}$ be the matrix of attention scores in a self-attention layer as in Definition S1.1. It follows that,
$\hat{\mathbf{X}}(\mathbf{X})=\mathbf{X}+\mathbf{A}(\mathbf{X})\mathbf{V}(\mathbf{X})=\mathbf{X}+\sigma\left(\frac{1}{\sqrt{d}}\,\mathbf{X}\mathbf{W}_{qk}\mathbf{X}^{T}\right)\mathbf{X}\mathbf{W}_{v}\,,$ (S11)
where $\mathbf{W}_{qk}=\mathbf{W}_{q}\mathbf{W}_{k}^{\top}$, and where we refer to a general layer in a Transformer model, omitting the subscript $l$. It follows that the entry $\alpha_{ij}=[\mathbf{A}]_{ij}$ of the attention score matrix is given by,
$\alpha_{ij}=\frac{\exp{(\mathbf{x}_{i}^{\top}\mathbf{W}_{qk}\mathbf{x}_{j})}}{\sum_{j}\exp{(\mathbf{x}_{i}^{\top}\mathbf{W}_{qk}\mathbf{x}_{j})}}=\frac{\exp{(\hat{\alpha}_{ij})}}{\mathcal{N}_{i}}\,,$ (S12)
where we neglect the scaling factor $1/\sqrt{d}$ for simplicity, every $\hat{\alpha}_{ij}$ is a coefficient given by $\hat{\alpha}_{ij}=\mathbf{x}_{i}^{\top}\mathbf{W}_{qk}\mathbf{x}_{j}$, and $\mathcal{N}_{i}$ is a normalization factor defined by the row-wise softmax function $\sigma(\cdot)$ and shared across all entries in the $i$-th row of $\mathbf{A}$,
$\mathcal{N}_{i}(\mathbf{X};\mathbf{W}_{qk})=\sum_{j}\exp{(\mathbf{x}_{i}^{\top}\mathbf{W}_{qk}\mathbf{x}_{j})}=\sum_{j}\exp{(\hat{\alpha}_{ij})}\,.$ (S13)
The coefficients $\{\hat{\alpha}_{ij}\}$ in the second term represent the projection of $\mathbf{x}_{i}$ onto $\text{span}\{\mathbf{X}\}$ (the subspace spanned by $\mathbf{X}=\{\mathbf{x}_{0}^{\top},\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{N}^{\top}\})$ in the transformed embedding space defined by $\mathbf{W}_{qk}$,
$\sum_{j}\hat{\alpha}_{ij}\,\mathbf{x}_{j}=\sum_{j}\mathbf{x}_{i}^{\top}\mathbf{W}_{qk}\mathbf{x}_{j}\,\mathbf{x}_{j}=\sum_{j}\langle\mathbf{x}_{i},\mathbf{W}_{qk}\mathbf{x}_{j}\rangle\,\mathbf{x}_{j}\,.$ (S14)
It follows from the monotonically increasing property of the soft-max function that the coefficients $\{\alpha_{ij}\}$ defined in Equation (S12) preserve the order of the coefficients $\{\hat{\alpha}_{ij}\}$ while exponentially suppressing negative and near-zero values,
$\alpha_{ij}<\alpha_{ij^{\prime}}\,\Leftrightarrow\,\hat{\alpha}_{ij}<\hat{\alpha}_{ij^{\prime}}\quad\forall i,j,j^{\prime}\,.$ (S15)
Therefore, the coefficients $\{\alpha_{ij}\}$ specify a convex combination that restricts the resulting vector in the convex hull $\text{Conv}(\mathbf{X})\subset\text{span}\{\mathbf{X}\}$, where the $i$-th row of $\mathbf{X}$ corresponding to the $i$-th token in the sequence is transformed into the row vector $\hat{\mathbf{x}}_{i}$ as follows,
$\hat{\mathbf{x}}^{\top}_{i}=\mathbf{x}^{\top}_{i}+\sum_{j}\alpha_{ij}\,\mathbf{x}_{j}\mathbf{W}_{v}\,,$ (S16)
thus concluding the proof. ∎
@@
Each element of the sum in Equation (S14) implicitly defines an operator in the subspace spanned by $\mathbf{X}=\{\mathbf{x}_{0}^{\top},\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{N}^{\top}\}$ given the transformed embedding space defined by $\mathbf{W}_{qk}$,
@@
thus performing the same operations as in Equation (S14).
S1.3. Proof of Proposition 2.2 and related remarks
@@
Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of $N$ tokens from a vocabulary of dimension $V$. Let $\mathcal{L}(U)$ be the negative log-likelihood of each token $t_{i}$, expressed as
$\mathcal{L}(U)=\sum_{i}\mathcal{L}(t_{i})=-\sum_{i}\log p(t_{i}\,|\,\{t_{j}:j\in\mathcal{C}_{i}\})\,,$ (S22)
where $\mathcal{C}_{i}\subset[0,1,\dots,N]$ is the set of indices defining the set of tokens $\{t_{j}\}$ of the conditional probability distribution. Let $\mathbf{U}=[\mathbf{t}_{0},\mathbf{t}_{1},\dots,\mathbf{t}_{N}]$ be the sequence of $N$ one-hot encoded tokens $\mathbf{t}_{i}\in\mathbb{R}^{V}$ associated with $U$, where $V$ is the dimension of the vocabulary. Let $\mathcal{L}(\mathbf{t}_{i})$ be the cross-entropy of the one-hot encoded token $\mathbf{t}_{i}$ and the estimated probability distribution $\sigma(\mathbf{z}_{i})\in\mathbb{R}^{V}$, as follows,
@@
$a_{l}(\mathbf{x}_{i}^{l-1})=\sum_{j\in C_{i}}\alpha^{l}_{ij}(w^{l})\,{\mathbf{x}^{l-1}_{j}}^{\top}\mathbf{W}^{l}_{v}\,.$ (S26)
@@
$\begin{split}\frac{\partial\mathcal{L}(\mathbf{t}_{i})}{\partial\alpha^{l}_{ij}}&=\frac{\partial\mathcal{L}(\mathbf{t}_{i})}{\partial\mathbf{z}_{i}}\,\frac{\partial\mathbf{z}_{i}}{\partial\mathbf{x}^{L}_{i}}\,\frac{\partial\mathbf{x}^{L}_{i}}{\partial\hat{\mathbf{x}}^{l}_{i}}\,\frac{\partial\hat{\mathbf{x}}^{l}_{i}}{\partial\alpha^{l}_{ij}}\\ &=(\mathbf{t}_{i}-\sigma(\mathbf{z}_{i}))^{\top}\mathbf{W}_{u}^{\top}\,\frac{\partial\mathbf{x}^{L}_{i}}{\partial\hat{\mathbf{x}}^{l}_{i}}\,{\mathbf{W}^{l}_{v}}^{\top}\sum_{j\in C_{i}}\mathbf{x}^{l-1}_{j}\,,\end{split}$ (S28)
@@
$\nabla_{w^{l}}\mathcal{L}(\mathbf{t}_{i})={\mathbf{\delta}^{l}_{i}}^{\top}\sum_{j\in C_{i}}\mathbf{x}^{l-1}_{j}\frac{\partial\alpha^{l}_{ij}}{\partial w^{l}}\,.$ (S32)
@@
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(\mathbf{t}_{i})={\mathbf{\delta}^{l}_{i}}^{\top}\sum_{j\in C_{i}}\mathbf{x}^{l-1}_{j}\mathbf{K}^{l-1}_{ij}\,,$ (S35)
@@
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(\mathbf{t}_{i})=\sum_{j\in C_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (S36)
@@
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(U)=\sum_{i}\sum_{j\in C_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\quad\forall\,\mathbf{W}_{qk}^{l},\,l\in[1,L]\,.$ (S37)
Equivalently, we can rewrite the double summation as follows,
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(U)=\sum_{i\in P_{j}}\sum_{j}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\quad\forall\,\mathbf{W}_{qk}^{l},\,l\in[1,L]\,,$ (S38)
where $P_{j}\subset[0,1,\dots,N]$ is the set of indices defining the set of tokens $\{t_{i}\}$ that are predicted by $t_{j}$, thus concluding the proof. ∎
@@
S1.4. Formal proofs for the connection between objective functions, directionality, and symmetry
@@
In this section, we provide a formal mathematical derivation of the informal Theorems 2.3 and 2.4 introduced in the main text. To do so, we provide a formal description and proof of the Theorems for the directionality and symmetric structures during autoregressive and bidirectional training, respectively with Theorem S1.12 and Theorem S1.14. We also introduce a series of related Propositions and Lemmas.
S1.4.1 Different implicit update of rows and columns

### v2

@@ 4.1. Evolution of symmetric and directional structures during learning
Figure 3: a) Evolution of symmetry score during training. We train a Bert-base-uncased model with 12 layers on the Wikipedia dataset [Foundation, 2022] in encoder-only (blue) and decoder-only (purple) mode, see legend. Shown are the median and the interquartile range. Shown are the median and the interquartile range. b) Same as in panel a for the median directionality score. c) Evolution of the median symmetry score across layers of the encoder-only (left) and decoder-only (right) models. Each layer is color-coded as shown on the legend. d) Same as panel c for the median directionality score.
To test the applicability of our result, we first train 12-layer transformer models in both encoder and decoder modes and quantify the median symmetry and directionality scores across epochs. At initialization, the symmetry and directionality score of the matrix $\mathbf{W}_{qk}$ at any layer is zero (see Definition 3.1 and Definition 3.2 and related Appendix S1.6 and S1.7). The incremental update of $\mathbf{W}_{qk}$ we described in the previous sections predicts that decoder-only models develop high-norm columns incrementally during training (see Theorem 2.3). Likewise, as symmetric weight updates are added to $\mathbf{W}_{qk}$ in encoder-only models, Theorem 2.4 predicts that symmetric structures emerge incrementally during training.
@@ 4.2. Enforcing symmetry at initialization improves the training of encoder-only models
Table 1: The final loss at the end of training and the speed-up for the 4, 12, and 24-layer models trained on the Jigsaw dataset [cjadams et al., 2017], Wikipedia [Foundation, 2022], and Red Pajama [Computer, 2023], with and without symmetry initialization (see Appendix S2.1). Speed-up (%) is calculated by subtracting the epoch at which the symmetrically initialized model reaches the non-symmetric model’s final loss from the total number of epochs, and then dividing by the total number of epochs. For example, a 50% speed-up means that the model with symmetric initialization achieves the final loss of the non-symmetric model in half the number of training epochs.
@@
24-layer model
Jigsaw 0.786
Jigsaw (+ symm) 0.739 14%
Wikipedia 0.192
Wikipedia (+ symm) 0.166 44%
Red Pajama 0.209
Red Pajama (+ symm) 0.189 34%
@@
We train 4-layer, 12-layer, and 24-layer encoder-only models, comparing two initialization strategies: initialize the self-attention matrices independently versus initializing the $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ in each self-attention layer to ensure that $\mathbf{W}_{qk}$ is symmetric (see Appendix S2). We report the results of our experiments in Table 1. We observe that enforcing symmetry at initialization leads to lower loss values at the end of training for most of the models. Importantly, symmetric initialization significantly accelerates convergence, reaching the final loss value faster than those with random initialization (up to ${}75\%$ faster for 4-layer models, ${}35\%$ faster for 12-layer models, and ${}44\%$ faster for 24-layer models; see also Figure S5a). Moreover, we observe that self-attention matrices initialized symmetrically lose symmetry during training but converge to higher symmetry levels than random initialization (Figure S5b) This symmetric initialization decreases the gap in symmetric scores between layers compared to random initialization (Figure S5c). When we tested symmetric initialization on vision Transformers, the improvements were not significant compared to the gains observed in language models (see Appendix S3 for a detailed analysis). These results highlight that embedding symmetry as an inductive bias across all Transformer layers can enhance training efficiency and model performance.
@@ Acknowledgments
MS was supported by an ETH Postdoctoral fellowship (reference nr. 23-2 FEL-042). PVA was supported by a UZH Postdoctoral fellowship (reference nr. K-76207-01-01). TS and PS were supported by DIZH Fellowship “Stability of self-organizing net fragments as inductive bias for next-generation deep learning” from the ZHAW. We thank Yassine Taoudi-Benchekroun, the Grewe lab, Alexander Nedergaard, Thomas Hofmann, and the Data Analytics lab for their helpful comments and insightful discussions.
@@
(Autoregressive and bidirectional objectives). Let $U=\{t_{1},\dots,t_{N}\}$ a sequence of tokens. The joint probability of $U$ is factorized autoregressively as follows,
@@
During autoregressive training, a model with a set of parameters $\mathcal{W}$ is optimized to minimize the following negative log-likelihood
$\mathcal{L}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(t_{i}\,|\,\{t_{j}\}:j<i)\,,$ (S7)
where the conditional probabilities are modeled with learnable parameters $\mathcal{W}$. During bidirectional training, a model with a set of parameters $\mathcal{W}$ is optimized to minimize the following negative log-likelihood
$\mathcal{L}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(t_{i}\,|\,\{t_{j}\}:j\neq i)\,.$ (S8)
@@
$\mathcal{L}(U;\mathcal{W})=-\sum_{i\in M}\log p_{\mathcal{W}}(t_{i}\,|\,\{t_{j}\}:j\notin M)\,.$ (S9)
S1.2. Related remarks to Section 2.1
@@
Each element of the sum in Equation (4) implicitly defines an operator in the subspace spanned by $\mathbf{X}=\{\mathbf{x}_{0}^{\top},\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{N}^{\top}\}$ given the transformed embedding space defined by $\mathbf{W}_{qk}$,
@@
thus performing the same operations as in Equation (4).
S1.3. Related remarks to Section 2.2
Let $\mathcal{V}=\{t_{1},\dots,t_{V}\}$ be a vocabulary of tokens, and let $D$ be a dataset of sequences $U_{u}\in\mathcal{V}^{N}$ of length $N$, where $D=\left\{U_{u}=(x^{u}_{1},\dots,x^{u}_{N})\right\}_{u=1}^{D}\subset\mathcal{V}^{N}$. For a single sequence $U_{u}$ with joint distribution $p(U_{u})$, the chain rule implies that for any permutation $\sigma$ of $\{1,\dots,N\}$, the joint probability can be written as,
$p(U_{u})=\prod_{t=1}^{N}p\big{(}x^{u}_{\sigma(t)}\mid x^{u}_{\sigma(1)},\dots,x^{u}_{\sigma(t-1)}\big{)}.$ (S15)
Given a set of conditional parents $\mathcal{D}_{i}\subset\{1,\dots,N\}\setminus\{i\}$, define the **local negative log-likelihood** as:
$\mathcal{L}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(x_{i}\mid x_{\mathcal{D}_{i}}),$ (S16)
where $\mathcal{W}$ denotes the model parameters. Let $G$ be a directed graph where each edge is $\mathcal{D}_{i}\to i$. We distinguish two cases:
  1. 1.
    If $G$ is a directed acyclic graph (DAG), then a topological ordering $\tau$ exists such that $\mathcal{D}_{\tau(i)}\subseteq\{\tau(1),\dots,\tau(i-1)\}$. In this case, the joint distribution $p_{\mathcal{W}}(x)=\prod_{i=1}^{N}p_{\mathcal{W}}(x_{\tau(i)}\mid x_{\mathcal{D}_{\tau(i)}})$ is properly normalized and the local negative log-likelihood in (S16) is equal to the true negative log-likelihood. For example, during autoregressive training, each token is conditioned only on its predecessors,
    $\mathcal{L}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(x_{i}\mid x_{1},\dots,x_{i-1})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(x_{i}\mid x_{<i}).$ (S17)
  2. 2.
    If $G$ contains directed cycles, the product $\prod_{i=1}^{N}p_{\mathcal{W}}(x_{i}\mid x_{\mathcal{D}_{i}})$ is not guaranteed to be normalized, and thus does not define a proper joint likelihood. In this case, one can instead minimize the following pseudo log-likelihood,
    ^$\hat{\mathcal{L}}(U;\mathcal{W})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(x_{i}\mid x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{N})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(x_{i}\mid x_{-i}),$ (S18)
    where $x_{-i}$ denotes all tokens in the sequence except $x_{i}$. Following Besag’s theorem Besag [1975], the maximum pseudo-likelihood estimator is consistent under standard regularity conditions, and (S18) remains compatible with stochastic-gradient optimization even though it is not the true negative log-likelihood Yang et al. [2019].
S1.4. Proof of Proposition 2.1 and related remarks
@@
Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of $N$ tokens. Let $\mathcal{L}(U;\mathcal{W})$ be the negative log-likelihood of each token $t_{i}$, expressed as
$\mathcal{L}(U;\mathcal{W})=\sum_{i=1}^{N}\mathcal{L}(t_{i})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(t_{i}\,|\,t_{\mathcal{D}_{i}})\,,$ (S19)
where $\mathcal{D}_{i}\subset[0,1,\dots,N]$ is the conditioning set, that is, the set of indices defining the set of tokens $\{t_{j}:j\in\mathcal{D}_{i}\}$ of the conditional probability distribution. Let $\mathbf{U}=[\mathbf{t}_{0},\mathbf{t}_{1},\dots,\mathbf{t}_{N}]$ be the sequence of $N$ one-hot encoded tokens $\mathbf{t}_{i}\in\mathbb{R}^{V}$ associated with $U$, where $V$ is the dimension of the vocabulary. Let $\mathcal{L}(\mathbf{t}_{i})$ be the cross-entropy of the one-hot encoded token $\mathbf{t}_{i}$ and the estimated probability distribution $\sigma(\mathbf{z}_{i})\in\mathbb{R}^{V}$, as follows,
@@
$a_{l}(\mathbf{x}_{i}^{l-1})=\sum_{j\in\mathcal{D}_{i}}\alpha^{l}_{ij}(w^{l})\,{\mathbf{x}^{l-1}_{j}}^{\top}\mathbf{W}^{l}_{v}\,.$ (S23)
@@
$\begin{split}\frac{\partial\mathcal{L}(\mathbf{t}_{i})}{\partial\alpha^{l}_{ij}}&=\frac{\partial\mathcal{L}(\mathbf{t}_{i})}{\partial\mathbf{z}_{i}}\,\frac{\partial\mathbf{z}_{i}}{\partial\mathbf{x}^{L}_{i}}\,\frac{\partial\mathbf{x}^{L}_{i}}{\partial\hat{\mathbf{x}}^{l}_{i}}\,\frac{\partial\hat{\mathbf{x}}^{l}_{i}}{\partial\alpha^{l}_{ij}}\\ &=(\mathbf{t}_{i}-\sigma(\mathbf{z}_{i}))^{\top}\mathbf{W}_{u}^{\top}\,\frac{\partial\mathbf{x}^{L}_{i}}{\partial\hat{\mathbf{x}}^{l}_{i}}\,{\mathbf{W}^{l}_{v}}^{\top}\sum_{j\in\mathcal{D}_{i}}\mathbf{x}^{l-1}_{j}\,,\end{split}$ (S25)
@@
$\nabla_{w^{l}}\mathcal{L}(\mathbf{t}_{i})={\mathbf{\delta}^{l}_{i}}^{\top}\sum_{j\in\mathcal{D}_{i}}\mathbf{x}^{l-1}_{j}\frac{\partial\alpha^{l}_{ij}}{\partial w^{l}}\,.$ (S29)
@@
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(\mathbf{t}_{i})={\mathbf{\delta}^{l}_{i}}^{\top}\sum_{j\in\mathcal{D}_{i}}\mathbf{x}^{l-1}_{j}\mathbf{K}^{l-1}_{ij}\,,$ (S32)
@@
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(\mathbf{t}_{i})=\sum_{j\in\mathcal{D}_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (S33)
@@
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(U)=\sum_{i}\sum_{j\in\mathcal{D}_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\quad\forall\,\mathbf{W}_{qk}^{l},\,l\in[1,L]\,.$ (S34)
We can rewrite the double summation either as,
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(U)=\sum_{i}\sum_{j\in\mathcal{C}_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\quad\forall\,\mathbf{W}_{qk}^{l},\,l\in[1,L]\,.$ (S35)
where $\mathcal{C}_{j}\subset[0,1,\dots,N]$ is the set of indices defining the set of tokens $\{t_{j}\}$ used to predict $t_{i}$, or equivalently as,
$\nabla_{\mathbf{W}_{qk}^{l}}\mathcal{L}(U)=\sum_{i\in\mathcal{P}_{j}}\sum_{j}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\quad\forall\,\mathbf{W}_{qk}^{l},\,l\in[1,L]\,,$ (S36)
where $\mathcal{P}_{j}\subset[0,1,\dots,N]$ is the set of indices defining the set of tokens $\{t_{i}\}$ that are predicted by $t_{j}$, thus concluding the proof. ∎
@@
S1.5. Formal proofs of Theorem 2.3 and 2.4: the connection between objective functions, directionality, and symmetry
@@
In this section, we provide formal proofs of Theorems 2.3 and 2.4, which characterize the directional and symmetric structures that emerge during autoregressive and bidirectional training, respectively. To this end, we introduce a series of intermediate Propositions and Lemmas, following these steps:
  1. 1.
    We begin by showing that when $t^{*}$ is used as a context token to predict other tokens, all predicted tokens contribute to the column space of $\mathbf{W}_{qk}$, while only $t^{*}$ contributes to the row space. Conversely, when $t^{*}$ is the token being predicted, only $t^{*}$ contributes to the column space, whereas all context tokens contribute to the row space (Section S1.5.1).
  2. 2.
    Next, we show that under reasonable assumptions about the statistical distribution of token embeddings, these structural properties imply that the expected norm of column updates exceeds that of row updates when $t^{*}$ is used as context. In contrast, row updates dominate when $t^{*}$ is the predicted token (Section S1.5.2).
  3. 3.
    We then show that in autoregressive training, the expected number of times a token $t^{*}$ appears as context can differ from the expected number of times it is predicted, since these two quantities are affected differently by statistical correlations between tokens. In contrast, under bidirectional training, the expected counts are always equal, regardless of token correlations (Section S1.5.3).
  4. 4.
    Finally, we combine the above results to prove our main theorems: (1) the weight updates to $\mathbf{W}_{qk}$ induce column dominance under autoregressive training (Section S1.5.4-S1.5.5), and (2) the weight updates to $\mathbf{W}_{qk}$ induce symmetry under bidirectional training (Section S1.5.6).
S1.5.1 Proof of Proposition 2.2